### v1

# Safe Text-to-Image Generation: Simply Sanitize the Prompt Embedding

Abstract
In recent years, text-to-image (T2I) generation models have made significant progress in generating high-quality images that align with text descriptions. However, these models also face the risk of unsafe generation, potentially producing harmful content that violates usage policies, such as explicit material. Existing safe generation methods typically focus on suppressing inappropriate content by erasing undesired concepts from visual representations, while neglecting to sanitize the textual representation. Although these methods help mitigate the risk of misuse to some extent, their robustness remains insufficient when dealing with adversarial attacks.
Given that semantic consistency between input text and output image is a fundamental requirement for T2I models, we identify that textual representations (i.e., prompt embeddings) are likely the primary source of unsafe generation. To this end, we propose a vision-agnostic safe generation framework, Embedding Sanitizer (ES), which focuses on erasing inappropriate concepts from prompt embeddings and uses the sanitized embeddings to guide the model for safe generation. ES is applied to the output of the text encoder as a plug-and-play module, enabling seamless integration with different T2I models as well as other safeguards. In addition, ES’s unique scoring mechanism assigns a score to each token in the prompt to indicate its potential harmfulness, and dynamically adjusts the sanitization intensity to balance defensive performance and generation quality. Through extensive evaluation on five prompt benchmarks, our approach achieves state-of-the-art robustness by sanitizing the source (prompt embedding) of unsafe generation compared to nine baseline methods. It significantly outperforms existing safeguards in terms of interpretability and controllability while maintaining generation quality.
Warnings: This paper contains depictions and discussions of explicit and other inappropriate content that some readers may find disturbing, distressing, and/or offensive.
1 Introduction
Diffusion models [11, 39], as the current state-of-the-art (SOTA) generative paradigm, drive the development of text-to-image (T2I) generation systems. These T2I models are guided by textual prompts and aim to generate realistic images that align with the provided text descriptions, finding widespread application in fields such as art creation [44] and graphic design [51]. Between 2022 and 2023 alone, T2I models generated over 15 billion images [6]. While T2I models demonstrate significant potential in terms of generation performance, they also face risks of unsafe generation. For instance, the BBC reported that AI-generated child sexual abuse materials were being widely sold over the internet, severely violating ethical and legal norms [5]. Therefore, ensuring that the content generated by T2I models adheres to usage policies [8, 25] is an urgent necessity.
To address this issue, many commercial T2I online services have implemented various safeguards. For instance, DALL-E [30] uses a Moderation API to detect input prompts that may violate usage policies. Stable Diffusion is equipped with a SafetyChecker [31] to filter out generated images containing explicit content. However, as passive defense mechanisms, input and output moderation typically offer coarse-grained protection, which may cause T2I models to overly intercept and refuse to generate any images. In contrast, safe generation methods [17, 7, 15] focus on erasing inappropriate concepts from the model’s internal representations. These methods aim to suppress the emergence of these concepts in the generated content, rather than indiscriminately blocking inputs or outputs. For instance, SafeGen [17] fine-tunes the model to remove visual representations linked to explicit concepts, ensuring that the generated content does not reflect these concepts. Such methods offer more fine-grained protection for T2I models, enabling them to produce high-quality, safe images even when given prompts with inappropriate concepts.
However, despite the fact that existing safe generation methods are capable of suppressing specific target concepts to certain extent, their robustness remains limited [43, 46]. Considering that semantic consistency between the input text and the generated image is a fundamental requirement for T2I models, which means that inappropriate prompts can be the primary source of unsafe generations by T2I models. Existing methods usually focus on fine-tuning the denoiser of a T2I model or modifying the denoising process to erase inappropriate concepts from the visual representation, neglecting the sanitization of inappropriate concepts in the prompt embeddings. This exposes their limitations, as these methods typically only suppress explicitly deleted concepts and lack sufficient generalization ability when handling synonyms or related concepts. Adversarial attacks on T2I models, such as SneakyPrompt [49], exploit this weakness by using synonym substitutions to bypass safeguards, leading to unsafe generation. Therefore, a natural question arises: Could we propose a robust safety generation framework by erasing inappropriate concepts at the source (i.e., prompt embeddings)?
Figure 1: ES operates by sanitizing prompt embeddings to generate safe images and providing interpretability to indicate harmful tokens.
Our work. In this paper, we provide an affirmative answer by introducing a novel, vision-agnostic safe generation framework, Embedding Sanitizer (ES). As shown in Figure 1, ES operates as a plug-and-play module applied after the text encoder, without necessitating modifications to any T2I model components. This design ensures compatibility, allowing ES to function independently or integrate seamlessly with other protective measures. ES focuses on erasing representations of inappropriate concepts from prompt embeddings, producing sanitized embeddings that guide the T2I model in generating appropriate content. A key feature of ES is its internal scoring network, which assigns a score to each token in an input prompt to gauge its potential harmfulness. For example, when processing the prompt ”A person in a birthday suit,” the token ”suit” receives a high score, indicating it as a significant contributor to inappropriate content generation. During the sanitization phase, ES dynamically adjusts the intensity of sanitization based on these scores, minimizing the impact on low-score (harmless) tokens while preserving strong sanitization effects on high-score (harmful) tokens. The scoring mechanism and adaptive sanitization function not only enhance the identification and management of inappropriate content but also provide interpretability and controllability in the generation process.
ES is a plug-and-play deep learning model that accesses only the text encoder during training, without involving other T2I model components. This design supports the use of a customized objective function for efficient end-to-end training. As a vision-agnostic safe generation framework, ES does not rely on image data for training. Instead, training begins with the selection of target concepts for erasure (e.g., nude) and their corresponding anchor concepts (e.g., dressed), providing explicit learning targets for sanitization. To augment training data, we use randomly synthesized prompts to pair toxic prompts $P_{t}$ (e.g., A nude person) with clean prompts $P_{c}$ (e.g., A dressed person). This approach produces abundant, diverse contextual samples, enabling ES to effectively learn to identify and sanitize inappropriate content across various scenarios. During training, a mean squared error loss is employed to encourage ES to align the embedding of the toxic prompt $P_{t}$ with that of the clean prompt $P_{c}$. Post-training, ES sanitizes input prompts in real-time to produce clean embeddings, ensuring that generated content adheres to usage policies.
Contributions. Our main contributions are as follows:
  • We propose and investigate a vision-agnostic safe generation method, Embedding Sanitizer (ES), designed as a plug-and-play module that sanitizes prompt embeddings in a controllable manner and provides interpretability for inappropriate elements within prompts. This modular design ensures orthogonal compatibility with existing safeguards, allowing ES to function independently or integrate seamlessly with other protective measures.
  • We introduce a synthetic data-driven training strategy that supplies ES with a nearly limitless variety of contextual training data. By specifying target and anchor concepts for removal, we reformulate the sanitization task as a transformation from target to anchor concepts, enabling efficient end-to-end training of ES.
  • We conduct a comprehensive evaluation of ES on five prompt datasets and compare it with nine latest safeguard baselines. Results demonstrate that ES achieves SOTA robustness while maintaining content quality, validating that prompt embedding sanitization can significantly enhance resilience to adversarial attacks. To facilitate future studies, we open-source our code in the following repository: https://anonymous.4open.science/r/Embedding-Sanitizer-166E/.
Ethical Considerations. For inappropriate content displayed in the manuscript, we apply masking measures to ensure that readers are not directly exposed to sensitive materials, thereby adhering to ethical standards for handling such content.
2 Background
2.1 Text-to-Image (T2I) Generation Models
Diffusion models have recently become a SOTA generative paradigm [10], employing a probabilistic framework to generate high-fidelity images from noisy representations. The training of diffusion models is divided into the diffusion process and the denoising process. In the diffusion process, the model will gradually add noise to the image so that it eventually approximates the standard Gaussian distribution. In the denoising process, the denoising network predicts and removes the added noise, progressively recovering the original image. The training objective of diffusion models is to minimize the difference between the predicted noise and the true noise at each timestep, typically expressed using a simple mean squared error (MSE) loss:
$\mathcal{L}_{\text{MSE}}=\mathbb{E}_{x,c,\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(x_{t},t,c\right)\right\|_{2}^{2}\right],$ (1)
where $x$ represents the original image, $x_{t}$ is the noisy image at timestep $t$, and $c$ represents the guiding condition, such as a prompt. The variable $\epsilon$ is the randomly sampled Gaussian noise, while $\epsilon_{\theta}$ indicates the predicted noise generated by the denoising network parameterized by $\theta$.
The ability of diffusion models to generate high-fidelity images largely depends on the efficient processing of data in the original high-dimensional pixel space. however, this also results in significant computational overhead. To address this issue, latent diffusion models (LDM)[33] propose a more efficient solution by executing the diffusion and denoising processes in a low-dimensional latent space, significantly reducing computational complexity. Stable Diffusion (SD) [40, 41, 42] is a typical representative of the LDM framework, which mainly consists of three core components:
  • Text Encoder. The text encoder utilizes a pre-trained CLIP model [29] to convert input text prompts into dense semantic embeddings. These embeddings guide the generation process through cross-attention mechanism, ensuring that the generated images align with the semantic meaning of the input text.
  • Image Decoder. The image decoder uses a pre-trained variational autoencoder (VAE) model [14] to map high-dimensional images into a low-dimensional latent space, improving computational efficiency. During the inference stage, the denoised latent representations generated by the diffusion process are decoded by the VAE into high-resolution images.
  • Denoiser. The denoising network uses the U-Net architecture [34], which takes noisy latent representations and text embeddings as input and progressively predicts and removes the noise at each time step, refining the latent representations into coherent image representations.
Compared to earlier generative models, Diffusion models offer a more stable training process, overcoming common issues such as mode collapse [2]. However, despite the breakthroughs in the generation performance of T2I models, challenges remain in terms of safety [52, 50]. Ensuring that the generated images comply with safety guidelines and preventing the production of inappropriate, harmful, or biased content are still pressing issues that need to be addressed.
2.2 Adversarial Attacks on T2I Models
With the support of large-scale image-text pair datasets such as LAION-5B [37], the performance of T2I models achieves significant progress. However, certain content within these datasets may contain sensitive information (such as adult content or hate symbols), which the models may inadvertently memorize during training, subsequently generating harmful images. Despite the integration of various safeguards in T2I models to avoid this issue, they remain vulnerable to adversarial attacks.
Adversarial attacks against T2I models aim to create adversarial prompts that bypass or evade the model’s safeguards, generating inappropriate images that do not meet expected standards. These attacks typically manipulate or modify the input prompts, introducing adversarial elements while preserving semantic similarity to the original prompts, thereby rendering the model’s safeguards ineffective. Based on the adversary’s knowledge of the target model, adversarial attacks can be categorized into white-box attacks and black-box attacks. In white-box attacks, the adversary has full access to the internal information of the target model, including its architecture, parameters, and gradient calculations. This allows the adversary to directly use the internal knowledge of the model to construct adversarial samples. For example, by utilizing gradient-based optimization techniques, the adversary can identify a pseudo token to compel the model to generate specific inappropriate content [26].
In contrast, black-box attacks are conducted without access to the internal structure or parameters of the model, making them more reflective of real-world threat scenarios. Adversaries typically rely on locally deployed surrogate models to optimize adversarial prompts, leveraging the transferability of prompts to attack the target model. For instance, SneakyPrompt [49] uses a Stable Diffusion model equipped with an output detector as a surrogate model and applies reinforcement learning algorithms to automatically search for substitutes for sensitive words in prompts. These substitutes are optimized using a reward function based on safety compliance and semantic fidelity, encouraging the generated adversarial prompts to bypass safeguards. Ring-A-Bell [43] utilizes a CLIP model to extract context-independent sensitive concept vectors and add them to the target prompts, and then uses a genetic algorithm to search for semantically similar adversarial prompts. Similarly, MMA-Diffusion[46] employs gradient-driven optimization and sensitive word regularization to construct adversarial prompts that contain no sensitive words while maintaining semantic similarity to the target prompt.
3 Related Work and Motivation
In this section, we review and summarize existing safeguards in the T2I model. We then summarize the properties that an ideal safe generation framework should possess. By analyzing these properties, we highlight the limitations of existing safe generation methods, which motivates us to propose a new defense framework that satisfies these properties.
3.1 Safeguards for T2I Models
To address adversarial attacks and mitigate abuse risks, T2I models integrate one or more safeguards to prevent unsafe content generation [31]. These safeguards can be broadly categorized into three types according to their defense goals: input moderation, output moderation, and safe generation.
Input and Output Moderation. Input moderation serves as the first line of defense by operating at the prompt input stage. Before processing the user’s prompt, this mechanism identifies and blocks malicious prompts containing sensitive terms to prevent the generation of inappropriate content. For example, commercial models like DALL·E 3[3] and Imagen[9] employ text classifiers (e.g., Moderation API [24]) to detect prompts that may violate usage policies and reject these inputs before they reach the model. Output moderation acts as the final defense at the image output stage. Once a T2I model generates an image, it is screened by a violation detector to identify any inappropriate content. Only compliant images are returned to the user. For example, Stable Diffusion employs a SafetyChecker [31] built on the CLIP architecture. It contains predefined sensitive text embeddings for 17 explicit concepts. It calculates the cosine similarity between the embedding of the generated image and the inappropriate concepts. If the similarity exceeds a threshold, the model outputs a black image as a substitute.
Safe Generation. Safe generation, also known as concept erasure [21], is a type of machine unlearning method [4] applied at the image generation stage. This method aims to erase inappropriate target concepts from the model’s internal representations, ensuring appropriate content output even when users provide malicious prompts. Negative Prompt [1] is a concept suppression mechanism integrated in Stable Diffusion that enables the user or the system to specify concepts that should not be generated in order to avoid generating inappropriate content. In the denoising process, the method suppresses the inappropriate content by introducing the embedding of inappropriate concepts to add noise to the intermediate denoising result. SLD [35] combines classifier-free guided text conditioning with inappropriate concepts to suppress the generation of these concepts similarly to Negative Prompt. ESD [7] fine-tunes the cross-attention or non-cross-attention layers of the denoiser on target concepts, embedding the functionality of Negative Prompt into the model parameters. Similarly, CA[15] fine-tunes either the full denoiser or only its cross-attention on pairs of target prompts and anchor images, aiming to link inappropriate target concepts with safe and appropriate images. Safe-CLIP[27] relearns a safe CLIP model through contrastive learning between text and image pairs, aiming to break associations between inappropriate textual and visual concepts. SafeGen[17] focuses on eliminating sexually explicit content by fine-tuning the self-attention layers of the denoiser in a text-independent manner, forcing the model to apply a mosaic effect to any generated explicit content.
Properties Analysis of Safe Generation. To comprehensively guard against potential evasion strategies, such as adversarial attacks, the safe generation framework for T2I models should meet the following core criteria [20]: (i) Effectiveness. The framework must effectively erase target concepts so that they do not appear in the generated image. (ii) Robustness. The framework should possess strong robustness, capable of suppressing synonymous concepts of the target and maintaining its effectiveness against adversarial attacks. (iii) Specificity. When generating non-target concepts, the output images should align semantically with the text descriptions, while maintaining a distribution consistent with the original model. (iv) Fidelity. After erasing target concepts, the quality of the generated images should not be significantly impacted, ensuring that the resulting images remain high-quality and usable.
Furthermore, to enhance defense performance and flexibility, an ideal safe generation framework should satisfy the following advanced criteria: (v) Interpretability. The framework should be able to reveal the contributions of tokens to the generation of inappropriate content, enabling T2I model providers to understand the extent of specific inputs’ impact on outputs, facilitating debugging and improvement. (vi) Compatibility. The framework should operate independently of existing defense systems, allowing for seamless integration and collaboration with various protective measures. (vii) Controllability. The framework should allow for control over the degree of erasure of target concepts during use, enabling T2I model providers to dynamically adjust the defense strength based on specific application needs.
3.2 Limitations of Existing Safeguards.
Despite significant progress in the safeguards of T2I models, existing mechanisms still exhibit notable limitations when dealing with complex adversarial attacks [13]. First, input moderation effectively intercepts prompts containing common sensitive terms but performs inadequately when handling synonyms, paraphrases, and adversarial prompts [47]. Second, output moderation is constrained by the detection scope and performance of content detectors [28], which may fail to detect attacks such as prompt dilution [43]. Furthermore, input and output moderation typically provide only coarse-grained protection, often resulting in the blind interception of inappropriate content, which causes T2I models to refuse to generate any content, negatively impacting user experience [45, 16].
In contrast, safe generation methods provide finer-grained control over the target concepts, allowing for the generation of safe and appropriate outputs even in the presence of prompts containing inappropriate concepts. This is because safe generation methods focus on eliminating or suppressing the generation of inappropriate concepts rather than completely blocking inputs or outputs, ensuring that users receive meaningful and compliant results when using the model. Although safe generation methods perform well in actively suppressing the generation of inappropriate concepts, they still lack robustness and are vulnerable to various adversarial prompts. Additionally, existing methods face further challenges in terms of interpretability, composability, and multi-concept erasure. Table I summarizes and compares the characteristics of influential existing safe generation methods. In terms of interpretability, existing methods primarily focus on suppressing the generation of specific concepts, making it difficult to pinpoint which tokens in the prompt trigger inappropriate outputs. This lack of interpretability hampers T2I service providers’ ability to effectively adjust and optimize defensive measures. Regarding compatibility, while SafeGen[17] and Safe-CLIP[27] claim seamless integration with other safeguards, SafeGen may encounter objective function conflicts with other fine-tuning-based methods such as ESD and AC. The secure text encoder built by Safe-CLIP can weaken the effectiveness of safe generation methods relying on embedding inappropriate concepts, such as Negative Prompt and SLD. In terms of controllability, existing methods often fail to dynamically adjust the defense intensity based on real-world application needs, making it difficult to balance defense performance and generation quality, which in turn limits the quality and diversity of generated outputs. These limitations motivate us to develop more robust safe generation frameworks to counteract the threats posed by increasingly sophisticated adversarial attacks.
[…]
Method Modified Component Property
Interpretability Combinability Controllability
NP
SLD
ESD U-Net
CA U-Net
Safe-CLIP Text-Encoder
SafeGen U-Net
ES (Ours)
4 Embedding Sanitizer
4.1 Threat Model
Following previous work [17], we define a threat model to describe the goals and capabilities of adversaries and T2I service providers, thereby clarifying the scope of our proposed defense framework.
Goals and Capabilities of Adversaries. The primary goal of adversaries is to bypass the safeguards in T2I models and induce the models to generate high-quality inappropriate or restricted content. We assume that adversaries have the capability to locally deploy an open-source T2I model, although this model fails to meet quality requirements. Adversaries can utilize the local model to create or collect any adversarial prompts, such as searching for adversarial text through optimization techniques. Additionally, we assume that adversaries can interact with the target model by attempting various prompt variations and can understand the security measures equipped within the target model, but they cannot directly access the internal parameter information of the target model.
Goals and Capabilities of T2I Service Providers. The primary goal of T2I service providers is to prevent unsafe generation and ensure the quality of the generated images. In terms of capabilities, T2I service providers have complete access to the parameters of the T2I model, enabling them to optimize additional modules or update model parameters based on gradient information. Furthermore, service providers can integrate various safeguards, such as Negative Prompt, aimed at enhancing the safety of the T2I model from a visual perspective.
4.2 Overview
Based on the analysis of the properties of safety generation methods, our goal is to develop a safe generation framework that embodies these properties, providing comprehensive defense for T2I models in complex real-world scenarios to ensure generated content adheres to usage policies.
Key Idea. A T2I model essentially functions as a text-driven image generator, with its primary objective being to generate high-quality images that align with the given prompt. Ensuring semantic consistency between the input text and the output image is a core requirement of the T2I model. For a well-trained, normal T2I model, it is challenging to generate an unsafe image from a clean prompt that contains no inappropriate descriptions. For instance, it is nearly impossible for a T2I model to produce an image containing nudity when prompted with ”A photo of a cute dog”, as this would clearly violate consistency requirements. In other words, any inappropriate content in the generated image likely originates from inappropriate descriptions within the input prompt, suggesting that a T2I model’s primary security risks may arise from the text encoder rather than other components.
Based on this insight, we introduce Embedding Sanitizer (ES), a vision-agnostic safe generation defense framework designed as an add-on plugin that operates on prompt embeddings, aiming to erase target concepts at the source (i.e., prompt embeddings). This approach to concept erasure provides ES with two practical advantages. First, most existing safe generation frameworks focus on removing target concepts from visual representations—typically by modifying the denoiser or denoising process—while ignoring the importance to remove toxic representations within text embeddings. This characteristic renders ES independent of other safe generation frameworks, providing it with compatibility and allowing seamless integration with various safeguards. Second, ES ’s training process only requires access to the text encoder, without needing access to the denoiser or image encoder. This design enables ES to function as a plug-and-play module that can be effortlessly transferred across different T2I models.
Next, we formalize the key idea of ES. Assume that the T2I model to be fortified is equipped with a text encoder denoted as $\mathcal{F}_{t}$. Given a toxic prompt $P_{t}$ containing inappropriate descriptions, such as A nude person, ES aims to transform this into a clean prompt $P_{c}$ with similar semantics, such as A dressed person. This can be represented as:
$\text{ES}(\mathcal{F}_{t}(P_{t}))\approx\mathcal{F}_{t}(P_{c}),$ (2)
where ES takes the toxic prompt embedding $\mathcal{F}_{t}(P_{t})$ as input and performs sanitization to bring the output embedding close to that of the clean prompt $P_{c}$.
Figure 2: Schematic diagram of the architecture of ES.
Overall Pipeline. The entire execution process of ES is illustrated in  Figure 2. At a high level, when a user inputs a prompt to query the T2I (text-to-image) model, the text encoder first uses a tokenizer and embedding lookup to encode each token in the text prompt into embeddings. The text encoder then further processes these embeddings through a self-attention mechanism to extract high-level feature representations, which are fed as input to ES. Finally, ES performs a sanitization process on the embeddings, aiming to erase inappropriate concepts and transform them into clean embeddings devoid of any unsafe elements.
At a microscopic level, the embeddings output by the text encoder are first processed by EmbeddingNet (E-Net), which is designed to capture toxic embeddings containing inappropriate concepts within the original suspicious embeddings. ScoreNet (S-Net) then takes the toxic embeddings as input, assigning a score to each token constrained within the range [0, 1] by a Sigmoid activation function. For toxic tokens, S-Net is inclined to predict a higher score to strongly erase the target concept. For clean tokens, S-Net tends to assign a lower score, aiming to minimize interference with the generation of non-target concepts. The scores predicted by S-Net are multiplied with the toxic embeddings predicted by E-Net, and the suspicious concepts undergo sanitization by subtracting the toxic components from the suspicious embeddings in a residual connection. The sanitized toxic embeddings can then serve as an approximate representation of clean embeddings, providing guidance for generating appropriate content.
It is worth noting that ES only sanitizes embeddings for positive prompts and does not modify embeddings for negative prompts. This preserves the text encoder’s ability to represent inappropriate concepts, thereby not interfering with the operation of negative prompt-based safe generation frameworks, such as SLD.
4.3 Architectural Design
Designed with modularity in mind, ES functions as an independent component and can seamlessly integrate with other safeguards. This modular approach enables ES to serve as a plug-and-play module compatible with various T2I systems that share a same text encoder. As shown in Figure 2, ES’s architecture comprises two main components: E-Net and S-Net, which collaboratively perform the crucial sanitization function. Below, we provide a detailed description of each component’s architectural design and explain how these designs confer interpretability and controllability to ES. The model hyperparameter configurations for these two components are summarized in Table II.
Table II: Model architecture parameters for ES.
Model Parameter Value
E-Net (Transformer Encoder) num_hidden_layers 6
num_attention_heads 8
hidden_size 768
intermediate_size 3072
S-Net (MLP) num_hidden_layers 2
input_size 768
intermediate_size 768
output_size 1
E-Net. The design of E-Net is based on the Transformer encoder architecture and includes six attention modules to efficiently capture potential inappropriate content in the input prompt. Each attention module employs an 8-head multi-head attention mechanism to enable E-Net to semantically analyze various parts of the input, allowing for a deeper understanding of tokens containing inappropriate concepts. E-Net’s role is to capture harmful concepts from the original suspicious embedding, producing a set of toxic embeddings that represent inappropriate content. Suppose $P_{s}$ is a suspicious prompt potentially containing inappropriate concepts, and it is encoded by the text encoder $\mathcal{F}_{t}$ to obtain the initial suspicious embedding, denoted as $\text{Emb}_{s}=\mathcal{F}_{t}(P_{s})$. E-Net, represented by $\mathcal{F}_{\theta}$, takes these embeddings as input and extracts identifiable toxic embeddings $\text{Emb}_{t}$, which can be formulated as:
$\text{Emb}_{t}=\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{s}))=\mathcal{F}_{\theta}(\text{Emb}_{s}).$ (3)
S-Net. S-Net is a multilayer perceptron (MLP) consisting of two fully connected layers, with its input dimension matching the output dimension of the embedding produced by E-Net. To constrain its output to [0, 1], a Sigmoid activation function is applied. The primary function of S-Net is to assign a score to each token in the input prompt, evaluating its potential harmfulness and indicating its contribution to the generation of inappropriate content, thereby providing ES with interpretability. Ideally, S-Net assigns high scores to toxic tokens, such as nude, enabling stronger erasure of toxic embeddings during the subsequent sanitization step. On the other hand, it assigns low scores to benign tokens, such as dog, minimizing disruption to the generation of non-toxic content. Represented by $\mathcal{F}_{\phi}$, S-Net takes the toxic embedding extracted by E-Net, $\text{Emb}_{t}$, as input and predicts a set of scores $\mathcal{S}$ for each token, expressed as:
$\mathcal{S}=\mathcal{F}_{\phi}(\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{s})))=\mathcal{F}_{\phi}(\text{Emb}_{t}).$ (4)
Embedding Sanitization. In the final sanitization step, the toxic embedding predicted by E-Net are multiplied by the scores predicted by S-Net, dynamically adjusting the sanitization strength according to each token’s score. This adjustment strategy provides two main advantages for ES ’s performance. First, it ensures that tokens identified as more toxic (i.e., with higher scores) undergo stronger sanitization, thereby enhancing the erasure of inappropriate content. Second, it preserves appropriate content by applying weaker sanitization to tokens identified as non-toxic (i.e., with lower scores). Next, ES sanitizes the original suspicious concept by subtracting (erasing) the toxic components from the suspicious embedding $\text{Emb}_{s}$ via residual connection to obtain a clean embedding $\text{Emb}_{c}$ that no longer contains inappropriate concepts. The sanitization process is formulated as:
$\text{Emb}_{c}$ $=\mathcal{F}_{t}(P_{s})-\alpha\cdot\mathcal{F}_{\phi}(\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{s}))\cdot\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{s}))$ (5)
$=\text{Emb}_{s}-\alpha\cdot\mathcal{S}\cdot\text{Emb}_{t}.$
It is worth noting that we introduce a hyperparameter $\alpha$ to control the strength of sanitization in the inference phase. When $\alpha$ approaches 0, sanitization is weak, and ES ’s effect is minimal, resulting in $\text{Emb}_{c}$ being close to the output of the text encoder $\text{Emb}_{s}$. Conversely, as $\alpha$ increases, sanitization strength is significantly enhanced, effectively erasing harmful components from $\text{Emb}_{s}$ and producing a cleaner embedding $\text{Emb}_{c}$. A smaller $\alpha$ reduces interference with the input prompt, preserving more of the original semantics but with weaker sanitization, while a larger $\alpha$ enhances sanitization at the cost of some semantic preservation.
Therefore, $\alpha$ provides ES with controllability, enabling it to adjust sanitization strength flexibly according to the safety requirements of specific application scenarios, balancing between retaining content relevance and effectively erasing inappropriate concepts.
Figure 3: Overview of the training process for ES.
4.4 Training Details
ES is essentially an end-to-end deep learning model that takes the original prompt embedding as input and outputs the sanitized clean embedding, which in turn guides the T2I model to generate appropriate content. Below, we introduce the training process of ES in detail, including training data preparation and objective function customization.
Training Data Preparation. To sanitize prompt embeddings, we first select an anchor concept corresponding to the target concept to be erased, providing ES with a clear direction for sanitization. This approach enables us to frame the embedding sanitization as a transfer problem toward the anchor concept. When the T2I model receives a prompt containing a target concept, ES redirects this concept toward the predefined anchor concept to achieve embedding sanitization. We design three approaches to construct $\langle$ target concept, anchor concept $\rangle$ pairs:
  • Antonyms of the target concept, e.g., $\langle$ illegal, legal $\rangle$;
  • Hypernyms of the target concept, e.g., $\langle$ blood, liquid $\rangle$;
  • A neutral placeholder, e.g., $\langle$ weapon, __\__ $\rangle$.
With the default settings we tend to use antonyms as anchor concepts to provide a more effective sanitization direction for the target concept.
Considering that the performance of deep learning models largely depends on vast amounts of training data, we design an automated method for synthesizing prompts to provide ES with an almost limitless supply of training samples. Specifically, we view the synthetic prompt $P$ as a sequence of $n$ tokens drawn from the vocabulary $\mathcal{V}$, represented as $P=\{\text{token}_{1},\text{token}_{2},\ldots,\text{token}_{n}\}$. During the data synthesis process, we first convert the target concept set $\mathcal{C}_{t}=\{c_{t}^{1},c_{t}^{2},\ldots,c_{t}^{m}\}$ containing $m$ concepts and the anchor concept set $\mathcal{C}_{a}=\{c_{a}^{1},c_{a}^{2},\ldots,c_{a}^{m}\}$ into their corresponding token IDs using a tokenizer. Next, within the maximum input length allowed by the text encoder (e.g., SD-v1.4 allows a maximum of 77 tokens), we randomly select a number of tokens to compose $P$, expressed as $P=\{(\text{token}_{1},\text{token}_{2},\ldots,\text{token}_{n})\mid\text{token}_{i}\in\mathcal{V},\forall i=1,2,\ldots,n,n\geq 3\}$. Here, the requirement that $n\geq 3$ ensures that $P$ contains at least three tokens: a starting token, a random token, and an ending token, which guarantees that $P$ complies with input specifications and is not empty. Finally, we select random positions excluding the first and last tokens to replace them with random concepts from $\mathcal{C}_{t}$, thereby forming a toxic prompt containing a random number of target concepts, denoted as $P_{t}$. We then replace the target concepts in $P_{t}$ with their corresponding anchor concepts to create the clean prompt $P_{c}$. This dataset construction strategy enables ES to learn how to transform harmful prompts $P_{t}$ into clean prompts $P_{c}$, ensuring that the generated images reflect appropriate content without losing contextual relevance.
The example in Figure 3 provides an intuitive illustration of a randomly synthesized $P_{t}$ and its corresponding $P_{c}$, where red tokens indicate target concepts and green tokens represent anchor concepts. Automated data generation strategy significantly enhances the generalization capability and adaptability of ES. First, synthetic data allows us to generate a vast number of prompts containing target concepts in various contexts, assisting ES in learning to identify and erase inappropriate content across diverse scenarios. Second, the generation of synthetic data enables precise control over the combinations of target concepts and anchor concepts, allowing ES to effectively adapt to different application requirements. Furthermore, generating training data in an automated manner can substantially reduce manual annotation costs and mitigate dependence on large-scale real-world inappropriate datasets, ensuring that ES can be trained broadly while safeguarding privacy and minimizing ethical controversies.
Initial Naive Implementation. In our initial naive implementation, we adopt a simple and intuitive objective function aimed at guiding ES to learn the sanitization functionality by minimizing the difference between the sanitized toxic embedding and their corresponding clean embedding. Given a toxic prompt $P_{t}$ and its corresponding clean prompt $P_{c}$, ES’s initial training objective is to transform the sanitized embedding of $P_{t}$ into an approximate embedding of $P_{c}$. We define the sanitized embedding $\hat{\text{Emb}_{t}}$ of $P_{t}$ and the clean embedding $\text{Emb}_{c}$ of $P_{c}$ using the following equations:
$\begin{cases}\hat{\text{Emb}_{t}}=\mathcal{F}_{t}(P_{t})-\mathcal{F}_{\phi}(\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{t})))\cdot\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{t})),\\ \text{Emb}_{c}=\mathcal{F}_{t}(P_{c}),\end{cases}$ (6)
where $\mathcal{F}_{t}$ denotes the text encoder, and $\mathcal{F}_{\theta}$ and $\mathcal{F}_{\phi}$ represent E-Net and S-Net, respectively. Thus, the initial objective function of ES is to minimize the following loss:
$\mathcal{L}_{init}=\min_{\theta,\phi}\|\hat{\text{Emb}_{t}}-\text{Emb}_{c}\|_{2}^{2},$ (7)
where $\mathcal{L}_{init}$ is the mean squared error (MSE) loss, which serves as the objective function penalizing the distance between $\hat{\text{Emb}_{t}}$ and $\text{Emb}_{c}$. Based on the construction of the synthetic data, this objective function encourages ES to approximate target concepts in the embedding toward the anchor concepts while preserving the original predictions of non-target concepts. This design effectively erases potential inappropriate content in the embedding while retaining the integrity of non-target concepts.
However, while the initial objective function Equation 7 provides ES with a learning direction to shift target concepts toward anchor concepts, equipping it with basic sanitization capability, challenges remain in generalizing to synonyms of target concepts. The core issue lies in the difficulty of exhaustively capturing all tokens related to a single concept when constructing synthetic data. For instance, when erasing a pornographic concept, we only define one relevant concept pair, such as $\langle$ nude, dressed $\rangle$. In other words, if a synonym of nude (such as ”bare” as shown in Figure 3) appears in the randomly generated toxic prompt $P_{t}$, the data synthesis strategy would retain it in the clean prompt $P_{c}$ rather than replace it with dressed, as it is not listed in the target concepts. This means that $\mathcal{L}_{init}$ fails to supervise the synonyms of target concepts, subsequently limiting ES’s generalization capability.
Final Improved Implementation. To address this challenge, we design an improved implementation for ES aimed at enhancing the generalization capability to synonyms of target concepts. In the initial implementation, the clean prompt $P_{c}$ consists mainly of random tokens and anchor tokens. We naively use the embedding of $P_{c}$ directly as the supervision target, overlooking the possibility that random tokens in $P_{c}$ may include synonyms of the target concepts. Therefore, considering that $P_{c}$ may still contain inappropriate concepts during its construction, our improved implementation not only sanitizes $P_{t}$ but also sanitizes $P_{c}$ to ensure that $P_{c}$ serves as a cleaner embedding for more accurately supervising ES’s learning. Similar to Equation 6, we define the sanitized embedding of $P_{t}$, denoted as $\hat{\text{Emb}_{t}}$, and the cleaner embedding of $P_{c}$, denoted as $\hat{\text{Emb}_{c}}$, using the following equations:
$\begin{cases}\hat{\text{Emb}_{t}}=\mathcal{F}_{t}(P_{t})-\mathcal{F}_{\phi}(\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{t})))\cdot\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{t})),\\ \hat{\text{Emb}_{c}}=\mathcal{F}_{t}(P_{c})-\beta\cdot\mathcal{F}_{\phi}(\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{c})))\cdot\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{c})),\end{cases}$ (8)
where $\beta$ is a hyperparameter that determines the extent of sanitization applied to $P_{c}$, constrained within the range of [0,1]. The improvement in ES’s generalization to synonymous concepts through the sanitization of $P_{c}$ is primarily driven by the semantic similarity between target concepts and their synonyms. This result arises from the higher cosine similarity between them in the embedding space, where their proximity is relatively small, causing synonymous concepts to typically be located among the neighbors around the target concept [12, 48]. By sanitizing $P_{c}$, this similarity is utilized to identify potential synonymous concept embeddings near the target concept’s embedding. Consequently, potential synonyms in $P_{c}$ can be effectively recognized and transformed into anchor concepts, providing ES with cleaner and more accurate supervision, thereby enhancing its generalization capability for synonymous concepts.
Figure 3 illustrates the final training process of ES. The text encoder of the T2I model first encodes the toxic prompt $P_{t}$ and the clean prompt $P_{c}$ into embedding representations, which are then fed into ES for sanitization. During training, ES shares the same parameters for sanitizing both $P_{t}$ and $P_{c}$. The goal of sanitizing the embedding of $P_{t}$ is to convert its target concepts into anchor concepts. In contrast, the goal of sanitizing the embedding of $P_{c}$ is to convert potential synonyms of target concepts into anchor concepts, thereby providing a cleaner supervision target. Based on the embedding representations in Equation 8, the final objective function is customized as follows:
$\mathcal{L}_{final}=\min_{\theta,\phi}\|\hat{\text{Emb}_{t}}-\hat{\text{Emb}_{c}}\|_{2}^{2},$ (9)
where $\mathcal{L}_{final}$ is the MSE loss that penalizes the distance between $\hat{\text{Emb}_{t}}$ and $\hat{\text{Emb}_{c}}$. This objective function encourages ES to bring potential target concepts and their synonyms in the embedding closer to anchor concepts, further enhancing ES’s robustness.
5 Evaluation and Analysis
In this section, we begin by outlining the evaluation setting, which includes datasets, safeguard baselines, and evaluation metrics. We then conduct extensive experiments to evaluate the performance of ES across four core dimensions: effectiveness, robustness, specificity, and fidelity. Finally, we explore the impact of hyperparameters on ES’s performance to demonstrate its controllability.
5.1 Evaluation Setup
Dataset. We consider five datasets for evaluating ES. These include a handcrafted prompt dataset, I2P [35], for effectiveness evaluation. Three adversarial prompt datasets, constructed from SneakyPrompt (SP) [49], Ring-A-Bell (RAB) [43], and MMA-Diffusion (MMA) [46], are used for robustness evaluation. Additionally, an image-text pair dataset, COCO-2017[18], is utilized for evaluate specificity and fidelity. We provide a description of the dataset in Appendix A.
Baselines. Considering that ES is a safe generation method, we prioritize comparisons with similar approaches. Specifically, we evaluate six safe generation baselines: NP[1], SLD[35], ESD[7], CA[15], Safe-CLIP[27], and SafeGen[17], along with three additional safeguard baselines: SD-v2.1[41], LG[19], and SC [22]. Unless otherwise specified, we follow prior research [17, 7, 35] by using Stable Diffusion (version 1.4) as the base model for all experiments. We provide a description of the configuration of the baseline method in Appendix B.
Training Setup. We default to using ES to erase the seven inappropriate concepts in I2P unless stated otherwise. The $\langle$ target concept, anchor concept $\rangle$ pairs used in the experiments are as follows:
$\langle$ nude, dressed $\rangle$, $\langle$ illegal, legal $\rangle$, $\langle$ violent, peaceful $\rangle$,
$\langle$ hateful, loving $\rangle$, $\langle$ harassed, respected $\rangle$,
$\langle$ harmful, helpful $\rangle$, $\langle$ shocking, soothing $\rangle$,
where we use antonyms of the seven inappropriate concepts as their anchor concepts. In training ES, we automatically construct training samples using a synthesis strategy based on the above (target concept, anchor concept) pairs, with a batch size of 32 and an AdamW optimizer with a learning rate of $5\times 10^{-4}$ for 10,000 optimization steps. In our experimental setup, the complete training process takes approximately 1 hour of GPU time, thanks to the lightweight architecture design shown in Table II, where ES has only 43.12M parameters, enabling efficient training and inference. Additionally, we set the hyperparameters of ES, $\alpha$ and $\beta$, to 3.0 and 0.9 by default to balance defense performance with generation quality.
Performance Metrics. To comprehensively evaluate ES ’s performance, we analyze it across four dimensions: effectiveness, robustness, specificity, and fidelity, introducing specific metrics for each:
  • Erasure Rate (ER): We use ER as a metric for effectiveness and robustness. ER is calculated by employing an external moderator to detect the presence of inappropriate content in generated images. ER is defined as the reduction of inappropriate concepts in generated images after applying a safeguard. Let $N$ denote the number of inappropriate concepts generated by the original T2I model without any safeguard, and $\hat{N}$ the number after applying the safeguard; then ER is given by $ER=1-\frac{\hat{N}}{N}$. With a higher ER indicating more effective erasure of inappropriate concepts. In line with SafeGEN and SLD, we use the image content detectors to identify inappropriate content, NudeNet[23] for the detection of explicit content and Q16[36] for the detection of other inappropriate content such as violence and hate material.
  • CLIP Score: We use the CLIP Score to evaluate specificity and it calculates cosine similarity between the given text embedding and the image embedding using the CLIP model[29]. For a benign prompt, a higher CLIP Score indicates that the generated image better matches the prompt description, meaning that the safeguard has a minimal effect on unrelated concepts.
  • Frechet Inception Distance (FID) score: We use FID score as a metric to assess fidelity and it is a common metric for evaluating the quality of generated images by measuring the distance between a set of generated images and a reference set. A lower FID score indicates higher semantic similarity between the generated and reference images, thus higher quality. We use the 5,000 real images from the validation set of COCO-2017[18] as the reference set.
Experimental Environment. We implement ES using Python 3.12.4 and Pytorch 2.3.1. All experiments can be conducted on servers running Ubuntu 24.04 LTS, equipped with an Nvidia GeForce RTX 4090 GPUs with 24 GB of VRAM.
Table III: Defensive performance of ES compared to nine baseline methods in terms of the erasure rate of inappropriate content generated from handcrafted crafted and adversarial prompts.
Dataset (Number of Inappropriate Contents) ER (%) $\uparrow$
SD-v2.1 LG SC NP SLD ESD CA Safe-CLIP SafeGen Ours
I2P (Handcrafted Prompt) Hate (91) -2.20 40.66 - 60.44 52.75 54.95 -2.20 49.45 - 70.33
Harassment (266) -5.26 35.34 - 59.40 48.50 42.86 6.77 33.83 - 75.94
Violence (319) 11.29 35.42 - 52.66 49.84 28.84 9.09 43.57 - 84.01
Self-harm (309) 2.91 33.66 - 58.90 59.87 42.39 6.15 46.60 - 84.14
Sexual (674) 33.98 34.27 65.28 48.96 45.25 72.70 16.02 33.23 43.76 93.03
Shocking (423) 2.84 39.72 - 51.06 46.10 40.90 6.38 43.50 - 85.11
Illegal activity (255) 7.06 38.82 - 64.31 52.55 46.67 6.27 45.88 - 80.78
Overall 7.23 36.84 - 56.53 50.69 47.04 6.93 42.29 - 81.91
Adversarial Prompt SP (581) 25.13 39.07 64.37 56.11 49.05 70.57 16.52 38.38 44.23 91.22
RAB (3596) 30.34 49.78 96.77 10.17 3.89 31.89 2.50 42.88 81.23 98.92
MMA (1738) 70.25 77.67 65.88 25.43 9.78 76.01 16.46 79.86 94.53 94.99
Overall 41.91 55.51 75.67 30.57 20.91 59.49 11.83 53.71 73.33 95.04
Figure 4: Comparison of the erasure rates for different exposed body parts between ES and safeguard baselines. SD-v1.4 generated a total of 3596 exposed body parts on the RAB dataset, while our method reduced this number to 39.
5.2 Effectiveness and Robustness Evaluation
We compare ES with nine baseline protection measures, including six safe generation baselines: NP, SLD, ESD, CA, Safe-CLIP, and SafeGen, as well as three additional protection baselines: SD-v2.1, LG, and SC. We evaluate the effectiveness of ES and the baselines in suppressing seven inappropriate concepts on the I2P dataset and evaluate their robustness against adversarial attacks using three adversarial prompt datasets: SP, RAB, and MMA. It is notable that I2P provides parameters such as seed and guidance scale for each prompt to ensure reproducibility of generated content. For fair comparison, we use the additional information provided by the I2P dataset to control image generation. For other prompt datasets, we set a different seed for each prompt to control the generated content, thus reducing errors caused by randomness and fairly reporting the erasure performance of each protective measure under the same random conditions.
As shown in Table III, we first report the total number of inappropriate content generated by the base model (SD-v1.4 without any safeguards) on each dataset. For example, the original SD-v1.4 generates 91 instances of hateful content using the I2P dataset. We then used the ER to demonstrate the ability of all methods to erase inappropriate concepts. The results show that ES achieves SOTA level of effectiveness, with the best erasure performance for all misclassifications in the I2P dataset. Notably, since SC and SafeGen are only effective for explicit content, we evaluate their performance only in that category. For erasure of sexual content in the I2P dataset, ES achieves the highest ER of 93.03%, while the lowest among the baselines is CA at 16.02% and the highest is ESD at 72.70%. In terms of the total erasure effect across all inappropriate categories, ES achieves the highest average ER of 81.91%, followed by NP with an average ER of 56.53%. This indicates that ES surpasses the current best baselines in effectiveness by approximately 40%.
Additionally, ES also achieves SOTA robustness, showing the best average defense performance against three types of adversarial attacks. In terms of overall erasure performance against adversarial prompts, SC and SafeGen demonstrate stronger defense capabilities compared to other baselines. This may be because SC, as an output moderation mechanism, directly detects and filters the generated content without relying on control over the generation process. For SafeGen, its text-agnostic design makes it less affected by synonym transformations of target concepts. In contrast, ES achieves the highest average ER of 95.04%, indicating excellent synonym generalization ability for target concepts and effective resistance to inappropriate handcrafted or adversarial prompts.
Figure 4 further demonstrates the defense effectiveness of the baseline methods and ES on the RAB adversarial prompt dataset. We use NudeNet[23] to detect exposed content in the generated images, which can identify various categories of exposure, such as female breasts (F_Breast) and male genitalia (M_Genitalia). It is noteworthy that the RAB attack generates more explicit content compared to I2P by adding extra pornographic semantics to the target prompt (the number of exposed content increases from 674 to 3596). The significant increase in explicit content enhances the output filter SC’s detection capability, resulting in SC achieving the best defense performance among the baselines under the RAB attack. In contrast, other baseline methods show weaker erasure effects in specific areas, such as Safe-CLIP’s limited effectiveness in erasing female breasts, and SafeGen’s poor erasure performance on the buttocks. However, ES effectively suppresses the generation of all exposed body parts, with an ER of 98.92% for total exposed areas, achieving SOTA performance. In addition, the visual comparison between ES and the baseline methods provided in Figure 5 further validates the effectiveness of ES.
Figure 5: Visual comparison of generated content between ES and six safe generation baselines. The first four rows show the generation results for toxic prompts, while the last two rows display the results for benign prompts.
5.3 Specificity and Fidelity Evaluation
We use CLIP scores and FID scores to evaluate the specificity and fidelity of ES, comparing it with the base model SD-v1.4 without any safeguards and six safe generation baselines. In the experiments, we use the COCO-2017[18] validation set as the real image set to support the calculation of FID scores. This validation set consists of 5,000 clean image-text pairs, where each image is annotated with a corresponding description. We use these annotations as generation prompts to produce 5,000 clean images and fix the same random seed for each prompt to ensure fair comparisons, making the generated image set consistent in content and quantity with the real image set for comprehensive quality evaluation.
Table IV reports the quantitative results of each method in terms of specificity and fidelity. The results show that NP and SLD introduce noise indiscriminately to target concepts even with clean prompts, leading to the worst FID scores and significantly compromising the fidelity of generated images. Although ESD achieves the best FID score, it has the lowest CLIP score, indicating that while ESD can produce high-fidelity images, it compromises specificity, reducing the semantic alignment between prompts and generated content. This observation is confirmed by the visual results in Figure 5. CA performs well in both specificity and fidelity, likely at the cost of effectiveness and robustness. SafeGen, in comparison, suppresses only concepts related to explicit content and not all seven inappropriate concepts, thus having a smaller impact on specificity and fidelity.
In contrast, ES has a negligible effect on specificity and fidelity, achieving CLPI and FID scores of 30.22 and 26.12, respectively, compared to the original SD-v1.4’s CLPI and FID scores of 31.31 and 25.25, indicating a compromise of about 3.7%. This impact is attributed to the default setting of a higher sanitization intensity ($\alpha=3.0$) in ES to enhance defensive performance, which moderately sacrifices generation quality. However, considering that ES improves effectiveness and robustness by approximately 25% to 40% over the best baseline, this slight quality trade-off is deemed acceptable. Figure 5 presents a visual comparison of clean generation content between ES and baseline methods, demonstrating that ES effectively preserves the original style and overall layout of images.
Table IV: Generation quality of ES on the COCO-2017 dataset compared to the baseline methods.
Safeguard CLIP Score $\uparrow$ FID Score $\downarrow$
COCO-2017
SD-v1.4 31.31 25.15
NP 30.24 30.63
SLD 30.60 29.67
ESD 29.25 24.38
CA 31.20 25.34
Safe-CLIP 30.84 26.49
SafeGen 31.26 25.58
Ours 30.22 26.12
5.4 Exploration on Hyperparameters
Figure 6: Performance of ES across different hyperparameters $\alpha$ (where $\beta=0.9$) and $\beta$ (where $\alpha=3.0$).
Figure 7: Examples illustrating the controllability of ES.
The performance of ES is influenced by two key hyperparameters, $\alpha$ and $\beta$. The hyperparameter $\alpha$ is applied during the inference stage to control the sanitization strength of the prompt embedding, while $\beta$ is used during the training stage to handle potentially inappropriate synonyms in clean prompts $P_{c}$ within training sample pairs. We explore the impact of $\alpha$ and $\beta$ on ES by adjusting their values. Considering the computational cost of extensive parameter comparisons, we choose to measure the ER on the adversarial prompt dataset SP to evaluate defense performance, and use the COCO-2017 dataset to measure CLIP and FID scores for assessing generation quality.
As shown in Figure 6 we set the evaluation range of $\alpha$ and $\beta$ to $[0,4.5]$ and $[0,0.9]$ respectively, observing how ES’s performance changes as these parameters vary. The results indicate that as $\alpha$ decreases and $\beta$ increases, the ER generally rises, while CLIP and FID scores tend to worsen. This suggests that smaller values of $\alpha$ significantly enhance the sanitization effect of the prompt embedding, and larger values of $\beta$ enable better handling of target concept synonyms during training, guiding ES to learn through cleaner supervision signals. This leads to better defense performance but may come at the cost of generation quality. Our final experiments indicate that choosing $\alpha=3.0$ and $\beta=0.9$ provides an optimal balance between defensive performance and generation quality.
Moreover, $\alpha$ can be adjusted as a system variable post-deployment to further balance defense performance and generation quality based on practical needs. In Figure 7, we illustrate the generation results under different $\alpha$ values to verify the controllability of ES. Here, $\alpha=0$ represents the output of the original model. The first and second rows of the figure show sensitive content, where decreasing $\alpha$ results in progressively more appropriate clothing on figures, indicating that $\alpha$ effectively controls the defense strength. The third and fourth rows show benign content, where further decreasing $\alpha$ causes slight layout changes, which might explain the minor drop in generation quality.
6 Discussion
In this section, we analyze the interpretability and composability of ES, and compare the computational overhead introduced by ES. Last, we discuss some future works.
Figure 8: Examples illustrating the interpretability of ES.
(a) Interpretability Analysis. During the inference process of ES, S-Net provides interpretability by assigning scores between 0 and 1 to each token in the prompt. These scores measure the potential harmfulness, identifying the sources of inappropriate content generation. As shown in Figure 8, we analyze the interpretability of ES through examples of prompts and highlight three key observations.
Firstly, ES effectively erases target concepts. As illustrated in the first row of Figure 8, ES assigns high scores of 0.99 and 0.61 to the tokens ”nude” and ”person,” respectively. This is because ES suppresses ”nude” as a target concept during training to prevent explicit content, with the high score of 0.99 indicating that ES successfully identifies ”nude” and applies strong sanitization measures to erase it. Additionally, the semantic meaning of ”nude” is propagated to related tokens through the attention mechanism of the text encoder. For instance, as ”nude” functions as an adjective describing ”person,” this may explain why ”person” receives a high score of 0.61 in the context of ”nude.” Secondly, ES effectively recognizes synonyms of target concepts. As shown in the second and third rows of Figure 8, although ”topless” and ”bare” are not explicitly set as target concepts during training, they are synonyms of ”nude,” and ES assigns them high scores of 0.85 and 0.40, respectively. This indicates that ES demonstrates good generalization to synonyms of target concepts, effectively sanitizing them even if they are not explicitly defined during training. Thirdly, ES does not affect non-target concepts. As shown in the last two rows of Figure 8, ES tends to assign scores close to 0 to benign concepts. For example, ”A fashionable person,” which is unrelated to explicit content, receives low scores for all tokens. Notably, while the same token ”person” receives a high score in the context of ”nude”, it receives a low score in the context of ”fashionable,” demonstrating that ES does not assign identical scores to the same token indiscriminately but rather adjusts scores contextually. Additionally, the scores close to 0 indicate that the sanitization strength of ES is dynamically adjusted to the minimum, which explains why in the visualized results, ES’s benign outputs closely resemble those of SD-v1.4.
Table V: Performance of ES when combined with different T2I models and baseline methods.
Safeguard ER (%) $\uparrow$ CLIP $\uparrow$ FID $\downarrow$
I2P (sexual) SP RAB MMA COCO-2017
Ours + SD-v1.4 93.03 91.22 98.92 94.99 30.22 26.12
Ours + SD-v1.1 94.95 92.59 99.38 95.97 29.71 29.45
Ours + SD-v1.2 93.02 94.66 99.24 95.05 29.85 29.03
Ours + SD-v1.3 92.87 90.87 98.74 94.59 30.13 26.31
Ours + SD-v1.5 91.39 92.59 98.58 94.64 31.44 24.71
Ours + NP 93.62 92.08 98.90 94.82 29.34 36.33
Ours + SLD 94.06 93.45 99.27 95.74 30.15 32.46
Ours + ESD 96.14 97.24 99.13 97.81 26.89 36.23
Ours + CA 90.94 90.53 98.63 93.95 30.05 25.79
Ours + Safe-CLIP 93.94 92.25 99.02 95.10 29.63 26.44
Ours + SafeGen 93.72 93.11 98.97 95.45 30.41 25.72
(b) Compatibility Analysis. The compatibility of ES is reflected in two aspects. On one hand, ES is compatible with different T2I models because its training process only requires access to the text encoder and does not depend on other components of the T2I model. This allows ES to be easily transferred to various T2I models that share the same text encoder structure, such as transitioning from SD-v1.1 to SD-v1.5. On the other hand, ES is also compatible with existing protective measures, as it functions as an auxiliary module that is orthogonal to methods based on fine-tuning or adjustments to the generation process. This enables ES to be seamlessly integrated with other protective measures, thereby further enhancing the safety performance of T2I models. Table V presents the compatibility results of ES, where Ours+SD-v1.4 serves as the baseline performance of ES. The results indicate that ES is applicable to different versions of T2I models and can cooperate effectively with most protective measures, achieving improved defense performance. It is worth noting that performance may be lower than the baseline in certain configurations. For instance, the FID score of Ours+NP is suboptimal, and the ER score of Ours+CA is lower, which may be attributed to the limited effectiveness of the combined method itself, thereby impacting the overall performance.
(c) Overhead Comparison. Considering that ES is a plug-and-play module, it may introduce additional computational overhead in the T2I generation process. To assess its computational burden, we compare the average time required to generate an image using ES and the baseline methods. Specifically, we generate 100 images in the same environment and calculate the average generation time for each method. As shown in Figure 9, ESD, CA, Safe-CLIP, and SafeGen are based on fine-tuned models, so they do not introduce additional computational overhead. In contrast, SLD[35], due to its complex noise-adding process, incurs a significantly higher computational cost, with an average time of approximately 4.52 seconds per image. In comparison, although ES is a plug-and-play module, its lightweight design (only 43.12M parameters in ES) ensures efficient inference, making the average generation time per image is nearly identical to the baseline model (SD-v1.4), without introducing extra computational burden. This demonstrates that ES maintains efficient computational performance while achieving SOTA-level defense capabilities, making it suitable for real-time applications in practical scenarios.
Figure 9: Comparison of computational overhead of ES and baseline methods.
(d) Limitations and Future Works. This study takes an important step toward achieving an ideal safety framework for T2I models. In practical deployments, flexibly adjusting the defense strength remains a challenging task. We design a residual skip-connection-based sanitization method that allows the defense strength to be adjusted freely by tuning the sanitization parameter $\alpha$. However, as a global hyperparameter, $\alpha$ uniformly controls the suppression effect across all target concepts, which may not meet the need for fine-grained control over different target concepts. Achieving this through more advanced architectural designs would be valuable for future work.
Moreover, we observe that ES essentially constructs a clean text embedding, as it operates as a vision-agnostic defense framework at the output of the text encoder. This implies that ES could potentially be applicable beyond T2I tasks, benefiting any task that relies on textual modalities. For instance, text-to-video generation tasks that use text embeddings to guide video creation [38], and text-to-image retrieval tasks that leverage text embeddings for image search [32], could all gain from ES’s protective capabilities. Future work could explore validating our defense technique across more task scenarios that utilize text embeddings.
7 Conclusion
In this paper, we investigate the risk of inappropriate generation in T2I models. To address this issue, we introduce Embedding Sanitizer (ES), a novel safe generation framework. ES successfully prevents potential inappropriate generation at the source by sanitizing the prompt embedding, thereby enhancing the safety and reliability of T2I model-generated content. Our framework not only effectively avoids generating inappropriate concepts but also provides improved interpretability and controllability. Moreover, ES’s modular design allows it to be compatible with existing safeguards, further enhancing the system’s flexibility. Through extensive comparisons across five datasets and with nine baseline methods, the results confirm that our approach achieves SOTA performance, effectively balancing robustness and fidelity.
Appendix A Dataset Descriptions
  • Inappropriate Image Prompt (I2P): I2P is a handcrafted prompt dataset collected from lexica.art, consisting of 4,703 text prompts with inappropriate descriptions, each associated with at least one of the following seven categories: hate, harassment, violence, self-harm, sexual content, shocking images, and illegal activity. It is also equipped with a seed and guide scale to ensure the reproducibility of the generated content. We use this dataset to evaluate ES ’s effectiveness in erasing these seven inappropriate concepts.
  • Adversarial Prompt Datasets: Considering that certain safeguard baselines (e.g., SafeGen) are only applicable for erasing explicit content we focus on generating adversarial prompt datasets related to explicit content. (i) Ring-A-Bell (RAB). Based on the default settings from the Ring-A-Bell’s source code, we filter 931 prompts related to sexual content from the I2P dataset and apply the Ring-A-Bell attack to generate adversarial prompts, forming the Ring-A-Bell dataset. (ii) SneakyPrompt (SP). Similarly, we apply the SneakyPrompt attack on prompts in I2P related to sexual content, creating a SneakyPrompt dataset with 931 adversarial prompts. (iii) MMA-Diffusion (MMA). Additionally, we use a dataset of 1000 adversarial prompts published by MMA-Diffusion, focused on inducing T2I models to generate explicit content. These datasets allow us to evaluate ES ’s robustness under adversarial attacks.
  • COCO-2017: COCO-2017 is a widely used image-text dataset providing diverse scenes, object combinations, and real annotations, covering 80 object classes and background contexts from the real world. Following prior research, we use the validation set of COCO-2017 to evaluate ES ’s specificity and fidelity.
Appendix B Baseline Descriptions
  • Safe Latent Diffusion (SLD): We use the official implementation of SLD, covering four safety levels—weak, medium, strong, and maximum. Following default settings, we set the safety level to medium and erase the following built-in target concepts: ”an image showing hate, harassment, violence, suffering, humiliation, harm, suicide, sexual, nudity, bodily fluids, blood, obscene gestures, illegal activity, drug use, theft, vandalism, weapons, child abuse, brutality, cruelty.”
  • Negative Prompt (NP): NP is a built-in feature of Stable Diffusion that allows the user to enter or the system to set a negative prompt to be used to suppress target concept generation, and we set the seven inappropriate concepts in I2P as negative system prompts.
  • Erased Stable Diffusion (ESD): ESD can be divided into ESD-x, which fine-tunes only the cross-attention layers, and ESD-u, which fine-tunes only the non-cross-attention layers. We use the ESD-u setting as it is stated in the original paper to perform better for erasing global concepts like explicit content. Following the default settings, we erase concepts like ”hate, harassment, violence, suffering, humiliation, harm, suicide, sexual, nudity, bodily fluids, blood” and train the model for 1000 epochs with a learning rate of $1\times 10^{-5}$.
  • Concept Ablation (CA): We adopt Model-based CA to erase seven inappropriate concepts in I2P and fine-tuned the model’s cross-attention layer 1000 times at a learning rate of $2\times 10^{-7}$ according to the source code settings.
  • Safe-CLIP: Safe-CLIP removes seven inappropriate concepts in the I2P dataset by fine-tuning the text encoder. We evaluate its performance by loading the safe text encoder provided by Safe-CLIP.
  • SafeGen: SafeGen focuses on erasing explicit concepts. We use the open-source parameters provided by SafeGen to assess its performance.
  • Stable Diffusion v2.1 (SD-v2.1): Compared to SD-v1.4, SD-v2.1 employs an external filter to clean the training data and is retrained based on this curated data.
  • Latent Guard (LG): LG is an input moderation method that recognizes inappropriate prompts by performing similarity comparisons in the embedding space. We evaluate the performance of LG using open source weights provided by the authors and using a threshold of 9.0131 set by default in the source code.
  • Safety Checker (SC): SC is an output moderation tool equipped with Stable Diffusion to filter explicit content. We use the official SD implementation to evaluate its performance.